### v1

# Examining the influence of anisotropy on the fundamental mode of nonradial oscillation in quark stars on a complete general relativistic scheme

Abstract
The anisotropic influence on the $f$-mode frequency of oscillations and dimensionless tidal deformability of strange quark matter are analyzed by employing the nonradial oscillation equations for the complete general relativity frame and tidal deformability equations, which are derived and modified from their standard form to introduce the anisotropic factor. The fluid in the compact star follows the MIT bag model with vector coupling. For the anisotropic function, we use a local anisotropy, which is regular along the whole star and is null both at the center and on the star’s surface. We show that the $f$-frequency of oscillation and dimensionless tidal deformability change considerably with the anisotropy. Finally, we investigate the correlation between the dimensionless tidal deformability of the GW170817170817 event with the anisotropy.
1 Introduction
Since the first detection of gravitational wave (GW) signals from a binary neutron star (NS) merger, known as GW170817 and reported by the LIGO-Virgo Collaboration (LVC) [1], numerous researchers have made significant efforts to explore the microphysics of NSs in more depth, providing stringent constraints on the equation of state (EOS) of such stars and significantly narrowing the range of possible theoretical models. In other words, the GW analysis provides a promising avenue for probing the behavior of ultra-dense matter under extreme conditions. This can be achieved by studying the EOS through the gravitational waves generated during the merger of two compact stars. Nevertheless, the GW170817 event is only one of the many and varied multi-messenger sources that can be observed using ground-based GW detectors [2, 3, 4, 5, 6, 7, 8, 9]. Furthermore, future space-based GW observatories like DECIGO are expected to open a new window for probing new physics related to highly deformed compact objects, such as NSs, hybrid stars, and magnetars, which are currently inaccessible to ground-based detectors [10].
Asteroseismology is a powerful theoretical tool for gaining a deeper understanding of the behavior of dense matter inside compact stars [11, 12, 13, 14]. Numerous studies have explored pulsation modes assuming various compositions of the isotropic fluid within compact stars, including deconfined quark matter [15], color-flavor-locked strange quark matter [16], hadronic matter [17, 18], scenarios with density discontinuities [19], and hadronic matter admixed with dark matter [20], among others. These studies consistently show that nonradial oscillation modes are highly sensitive to the microphysical characteristics of the fluid within compact stars, leading to significant variations in their behavior. It is also worth mentioning that third-generation instruments are expected to be more sensitive to continuous GW signals from, e.g., $r$-mode or $f$-mode instabilities, from NS mountains [21, 22] or due to resonances in compact binary systems [23].
One of the most widely accepted assumptions in the study of compact stars is that the dense fluid within these objects is isotropic. Nonetheless, theoretical evidence suggests that various phenomena may induce anisotropy in fluids under extreme conditions [24, 25, 26], see also the review articles [27, 28] for extensive discussions on anisotropic pressure in self-gravitating systems. In addition, it has been shown in recent years that the inclusion of anisotropy in relativistic star systems leads to theoretical predictions that are consistent with several observational mass-radius measurements and tidal deformability constraints [29, 30, 31, 32]. Motivated by these studies, in the present work, we investigate the impact of anisotropic pressure on the $f$-mode (fundamental mode) oscillation frequency and tidal deformability of strange stars by employing the MIT bag model with vector coupling. For this purpose, we derive the nonradial oscillation equations and the regularity conditions for the perturbation variables near the center within the framework of general relativity for anisotropic matter. These equations, together with the deformability equations, are then integrated to assess the influence of anisotropy on the fundamental oscillation modes and stellar deformability.
It is important to note that, in a similar context, the influence of anisotropy on fundamental oscillation modes was previously analyzed in [33, 34], where the authors derived the nonradial oscillation equations and presented the corresponding regularity conditions near the center. However, the equations and regularity conditions derived in our work differ from those presented in [33].
This manuscript is structured as follows: Section 2 presents the hydrostatic equilibrium equations and the nonradial oscillations equations, including the anisotropic factor. Section 3 shows the EOS and the anisotropic profile equation, and in Section 4 we report the results in the analysis of effects of the anisotropy on the equilibrium, frequency of oscillations, and tidal deformability. In Section 5, we conclude. Finally, in Appendix A, the perturbation variables and the general form of the linearized perturbation equations for the anisotropic case are presented. In Appendix B, the perturbation functions explicitly expanded at the star center are presented, which are important to guarantee finite and numerically stable solutions. Throughout this article, we adopt geometric units by setting $G=1=c$ to simplify the equations and facilitate numerical calculations.
2 General relativistic equations
2.1 The static equilibrium equations
For completeness, we start by writing the Einstein field equation in the presence of anisotropic matter:
$G_{\mu\nu}=8\pi T_{\mu\nu},$ (2.1)
with the Greek indices $\mu,\nu,$ etc., run from 0 to $3$. $G_{\mu\nu}$ represents the Einstein tensor and $T_{\mu\nu}$ stands the stress-energy tensor that is given by:
$T_{\mu\nu}=\left(p_{r}+\sigma\right)g_{\mu\nu}+(\rho+p_{r}+\sigma)u_{\mu}u_{\nu}-\sigma k_{\mu}k_{\nu}.$ (2.2)
The variables $p_{r}$, $\rho$, and $\sigma$ depict, respectively, the radial pressure, the energy density, and the anisotropic factor. $g_{\mu\nu}$, $u_{\mu}$, and $k_{\mu}$ represent the spacetime metric tensor, the fluid’s $4$-velocity, unit radial vector, respectively. In addition, the $4$-velocity and the unit radial vector follow the equalities:
$u_{\mu}u^{\mu}=-1,\quad k_{\mu}k^{\mu}=1,\quad{\rm and}\quad k_{\mu}u^{\mu}=0.$ (2.3)
To analyze the static equilibrium configuration of spherically symmetric static compact stars, we consider the line element in the Schwarzschild-like coordinates ($t,r,\theta,\phi$) as follows:
$ds^{2}=-e^{2\Psi}dt^{2}+e^{2\Lambda}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (2.4)
where $\Psi=\Psi(r)$ and $\Lambda=\Lambda(r)$ are functions of the radial coordinate $r$ only.
Replacing the energy-momentum tensor (2.2) and the metric (2.4) on the Einstein field equation, we derive the hydrostatic stellar structure equations:
$\frac{dp_{r}}{dr}=-\frac{\rho m}{r^{2}}\left[1+\frac{p_{r}}{\rho}\right]\left[1+\frac{4\pi r^{3}p_{r}}{m}\right]e^{2\Lambda}+\frac{2\sigma}{r},$ (2.5)
$\frac{dm}{dr}=4\pi r^{2}\rho,$ (2.6)
$\frac{d\Psi}{dr}=-\frac{1}{p_{r}+\rho}\frac{dp_{r}}{dr}+\frac{2\sigma}{r(p_{r}+\rho)},$ (2.7)
with
$e^{2\Lambda}=\left(1-\frac{2m}{r}\right)^{-1}.$ (2.8)
As usual, the parameter $m$ represents the mass within the sphere radius $r$. Eq. (2.5) is known as the hydrostatic equilibrium equation, it is also known as the Tolman-Oppenheimer-Volkoff equation [35, 36], modified from its standard form to include the anisotropic factor $\sigma$, see Ref. [37].
Eqs. (2.5)-(2.7) are integrated from the center ($r=0$) to the surface of the star ($r=R$). This solution starts at the center of the star $r=0$, where:
$m(0)=0,\quad\rho(0)=\rho_{c},\quad p_{r}(0)=p_{rc},\quad\sigma(0)=0,\quad{\rm and}\quad\Psi(0)=\Psi_{c},$ (2.9)
and the star’s surface is determined by $p_{r}(r=R)=0$. Moreover, the interior line element connects smoothly with the exterior Schwarzschild vacuum metric at the star’s surface. Thus, the inner and outer potential metric functions are connected by the relation:
$e^{2\Psi(R)}=e^{-2\Lambda(R)}=1-\frac{2M}{R},$ (2.10)
with $M$ representing the total mass of the star. Eq. (2.10) depicts the boundary condition of $\Psi(R)$ at the star’s surface.
2.2 Nonradial perturbation equations
The study of the nonradial oscillations is investigated by decomposing the perturbed metric into a background metric $g^{(0)}_{\mu\nu}$, whose components are found in Eq. (2.4), plus the metric perturbation $h_{\mu\nu}$. This decomposition can be placed into the form:
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+h_{\mu\nu}.$ (2.11)
In pulsating compact stellar configurations, the fluid spacetime dynamics are described by the perturbed Einstein field equation and conservation of the energy-momentum tensor
$\delta G^{\mu}_{\varphi}=8\pi\delta T^{\mu}_{\varphi},$ (2.12)
$\delta\left(\nabla_{\mu}T^{\mu}_{\varphi}\right)=0,$ (2.13)
with
$\delta G^{\mu}_{\varphi}=g^{(0)\mu\beta}\delta R_{\varphi\beta}-\frac{1}{2}\delta_{\varphi}^{\mu}\left(g^{(0)\alpha\beta}\delta R_{\alpha\beta}-h^{\alpha\beta}R^{(0)}_{\alpha\beta}\right)-\frac{1}{2}g^{(0)\mu\beta}h_{\varphi\beta}R^{(0)}-h^{\mu\beta}G^{(0)}_{\varphi\beta}$, (2.14)
$\delta T^{\mu}_{\varphi}=\left(\delta\rho+\delta p_{r}+\delta\sigma\right)u^{\mu}u_{\varphi}+\left(\rho+p_{r}+\sigma\right)\delta u^{\mu}u_{\varphi}+\left(\rho+p_{r}+\sigma\right)u^{\mu}\delta u_{\varphi}+\delta p_{r}g^{(0)}_{\nu\varphi}g^{(0)\mu\nu}$
$+\delta\sigma g^{(0)}_{\nu\varphi}g^{(0)\mu\nu}+(p_{r}+\sigma)h_{\nu\varphi}g^{(0)\mu\nu}-(p_{r}+\sigma)g^{(0)}_{\nu\varphi}h^{\mu\nu}-\delta\sigma g^{(0)}_{\nu\varphi}k^{\mu}k^{\nu}-\sigma h_{\varphi\nu}k^{\mu}k^{\nu}-\sigma g^{(0)}_{\varphi\nu}\delta k^{\mu}k^{\nu}$
$-\sigma g^{(0)}_{\varphi\nu}k^{\mu}\delta k^{\nu}.$ (2.15)
To investigate the effects of anisotropy on the fluid pulsation mode emitted by compact stars, following [38] and adopting the Regge and Wheeler [39] gauge, the metric perturbation $h_{\mu\nu}$ for a given even-parity spherical harmonic function $Y_{\ell m}(\theta,\phi)$ is given by
$h_{\mu\nu}=\begin{pmatrix}He^{2\Psi}&H_{1}&0&0\\ H_{1}&He^{2\Lambda}&0&0\\ 0&0&Kr^{2}&0\\ 0&0&0&Kr^{2}\sin^{2}\theta\end{pmatrix}Y_{\ell m},$ (2.16)
with $H$, $H_{1}$, and $K$ depending on $t$ and $r$. Through the non-zero components of the perturbed Einstein field equation (2.12) and the perturbed components of the conservation of the energy-momentum tensor ($\delta\left(\nabla_{\nu}T^{\nu}_{1}\right)=0$ and $\delta\left(\nabla_{\nu}T^{\nu}_{2}\right)=0$), the nonradial oscillation equations for a general anisotropic profile can be obtained; see Eqs. (A.14)-(A.20). Since we concentrate our attention on normal modes, following Ref. [40, 41], we employ $H=r^{\ell}{\tilde{H}}e^{i\omega t}$, $H_{1}=i\omega r^{\ell+1}{\tilde{H}}_{1}e^{i\omega t}$, $K=r^{\ell}{\tilde{K}}e^{i\omega t}$, $W=r^{\ell+1}{\tilde{W}}e^{i\omega t}$, and $V=r^{\ell}{\tilde{V}}e^{i\omega t}$; where the functions $\tilde{H}$}, ${\tilde{H}}_{1}$, $\tilde{K}$}, $\tilde{W}$}, and $\tilde{V}$} depend on $r$ and $\omega$ represents the oscillation eigenfrequency. By using these five equalities and introducing the function $\tilde{X}$} defined by
${\tilde{X}}=\left(1+\frac{\partial\sigma}{\partial p_{r}}\right)^{-1}\left[\omega^{2}(\rho+p_{r}+\sigma){\tilde{V}}e^{-\Psi}-\left(\rho+p_{r}\right)\frac{{\tilde{H}}}{2}e^{\Psi}+e^{\Psi+2\Lambda}{\tilde{H}}\frac{\partial\sigma}{\partial g_{11}}\right]-\frac{p_{r}^{\prime}{\tilde{W}}}{re^{\Lambda-\Psi}},$ (2.17)
equations (A.14)-(A.20) yield
${\tilde{H}}_{1}^{\prime}=\frac{e^{2\Lambda}}{r}\left[{\tilde{H}}+{\tilde{K}}+16\pi(\rho+p_{r}+\sigma){\tilde{V}}\right]-\frac{{\tilde{H}}_{1}}{r}\left[\ell+1+\frac{2me^{2\Lambda}}{r}+4\pi r^{2}(p_{r}-\rho)e^{2\Lambda}\right],$ (2.18)
${\tilde{K}}^{\prime}=\frac{{\tilde{H}}}{r}+\frac{\ell(\ell+1)}{2r}{\tilde{H}}_{1}-\frac{{\tilde{K}}}{r}\left(\ell+1-\Psi^{\prime}r\right)+8\pi\left(\rho+P_{r}\right)\frac{e^{\Lambda}{\tilde{W}}}{r},$ (2.19)
${\tilde{W}}^{\prime}=-(\ell+1)\frac{{\tilde{W}}}{r}+re^{\Lambda}\left[\frac{e^{-\Psi}{\tilde{X}}}{p_{r}+\rho+\sigma}\frac{d\rho}{dp_{r}}-\ell(\ell+1)\frac{{\tilde{V}}}{r^{2}}-\frac{{\tilde{H}}}{2}-{\tilde{K}}\right$], (2.20)
[…]
$\left.\left.+\rho\right)\left[\Psi^{\prime}-\frac{1}{r}\right]-\frac{2}{r}e^{\Psi+2\Lambda}\frac{\partial\sigma}{\partial g_{11}}\right]+{\tilde{K}}\left[(p_{r}+\rho)\frac{e^{\Psi}}{2}\left[\frac{1}{r}-3\Psi^{\prime}\right]+\sigma e^{\Psi}\left[\frac{3}{r}-\Psi^{\prime}\right]\right]-{\tilde{X}}\left[\frac{\ell}{r}\right$.
$\left.+\frac{d\rho}{dp_{r}}\left[\frac{\sigma\Psi^{\prime}}{p_{r}+\rho+\sigma}+\frac{2\sigma}{r(p_{r}+\rho+\sigma)}-\frac{2}{r}\frac{\partial\sigma}{\partial\rho}\right]\right]+{\tilde{V}}\frac{e^{\Psi}}{r^{2}}\ell(\ell+1)\left[-(p_{r}+\rho)\Psi^{\prime}+\frac{2\sigma}{r}\right],$ (2.21)
${\tilde{H}}\left[3m+\frac{1}{2}(\ell+2)(\ell-1)r+4\pi p_{r}r^{3}\right]={\tilde{H}}_{1}\left[\omega^{2}r^{3}e^{-2(\Psi+\Lambda)}-\frac{\ell(\ell+1)}{2}\left(m+4\pi r^{3}p_{r}\right)\right]$
$+{\tilde{K}}\left[\frac{1}{2}(\ell+2)(\ell-1)r-\omega^{2}r^{3}e^{-2\Psi}-\frac{e^{2\Lambda}}{r}\left(m+4\pi r^{3}p_{r}\right)\left(3m-r+4\pi r^{3}p_{r}\right)\right]-8\pi e^{-\Psi}{\tilde{X}}r^{3}$
$-16\pi e^{-\Lambda}\sigma{\tilde{W}}r.$ (2.22)
Here, the anisotropic profile employed depends on the fluid and spacetime variables of the form $\sigma=\sigma(p_{r},g_{11})$ with $p_{r}=p_{r}(\rho)$; see, e.g., [42, 43]. Eqs. (2.17)-(2.22) can be reduced to the fourth-order system of linear equations presented in [41] by taking ${\tilde{H}}\rightarrow-{H}$, ${\tilde{H}}_{1}\rightarrow-{H}_{1}$, ${\tilde{K}}\rightarrow-{K}$, and $\sigma=0$; review also [44]. In addition, Eqs. (2.20) and (2.21) differ from the respective equations derived in Ref. [33] in all terms where the anisotropic factor appears.
To solve Eqs. (2.17)-(2.22), some conditions at the center ($r=0$) and at the surface of the star ($r=R$) are required. In the center of the star, as is realized in the isotropic case [40, 41], regularity conditions of the perturbative variables are imposed. These conditions are found by expanding the fluid variables and the spacetime perturbations by Taylor power series near $r=0$. These expansions, together with the perturbation equations (2.17)-(2.22) reduced through these expansions, are shown in Appendix B. In turn, outside the stellar structure configuration $r\geq R$, the fluid variables $p_{r},\rho$, and $\sigma$ and the fluid perturbation quantities $W$ and $V$ are zero and $m=M$. In addition, such as is realized in [19], we replace the variables ${\tilde{H}}_{1}$ and $\tilde{K}$} with the new ones $Z$ and $dZ/dr^{*}$ through the equalities:
$\hskip-28.45274pt{\tilde{K}}=\frac{n(n+1)r^{2}+3nMr+6M^{2}}{r^{\ell+2}(nr+3M)}Z+\frac{1}{r^{\ell+2}}\frac{dZ}{dr^{*}},$ (2.23)
$\hskip-28.45274pt{\tilde{H}}_{1}=\frac{nr^{2}-3nMr-3M^{2}}{r^{\ell+1}(r-2M)(nr+3M)}Z+\frac{r^{1-\ell}}{(r-2M)}\frac{dZ}{dr^{*}},$ (2.24)
giving rise to the Zerilli equation [45, 46, 47]. In Eqs. (2.23) and (2.24), $n=(\ell-1)(\ell+2)/2$ and $r^{*}$ is the so-called “tortoise" coordinate, which is related to the radial coordinate $r$ through the relation:
$r^{*}=r+2M\ln\left(\frac{r}{2M}-1\right).$ (2.25)
At the star’s surface ($r=R$), the Lagrangian perturbation $\Delta p_{r}$ must vanish, i.e.,
$\Delta p_{r}=-r^{\ell}e^{-\Psi}{\tilde{X}}=0,$ (2.26)
thus implying ${\tilde{X}}(r=R)=0$. Each mode solution is uniquely associated with a specific ($\ell,\omega$) pair; however, not every combination of ($\ell,\omega$) corresponds to a valid mode solution, since the spectrum is discrete.
2.3 Tidal deformability
The investigation of tidal Love numbers has emerged as a key aspect in understanding binary systems consisting of compact stars. In these systems, the gravitational pull exerted by one star induces tidal distortions in its companion. Such deformations, arising from external tidal forces, are characterized by the dimensionless tidal deformability parameter $\Lambda$, which is defined as
$\Lambda=\frac{2k_{2}}{3C^{5}}.$ (2.27)
In this equation, $C=M/R$ represents the stellar compactness, and $k_{2}$ is the quadripolar Love number, which can be written as
${k_{2}}=$ $\ \frac{8C^{5}}{5}(1-2C)^{2}[2+C(y_{R}-1)-y_{R}]\times\{2C[6-3y_{R}+3C(5y_{R}-8)]+4C^{3}[13-11y_{R}$
$+C(3y_{R}-2)+2C^{2}(1+y_{R})]+3(1-2C^{2})\times[2-y_{R}+2(y_{R}-1)]\ln(1-2C)\}^{-1},$ (2.28)
where $y_{R}=y(r=R)$. The function $y(r)$ follows the Riccati-type differential equation
$y^{\prime}r+y^{2}+y\left(K_{0}r-1\right)+K_{1}r^{2}=0,$ (2.29)
with coefficients
$K_{0}=$ $\ \frac{2m}{r^{2}}e^{2\Lambda}+4\pi e^{2\Lambda}(p_{r}-\rho)r+\frac{2}{r},$ (2.30)
$K_{1}=$ $\ 4\pi e^{2\Lambda}\left[4\rho+8p_{r}+4\sigma+\frac{p_{r}+\rho}{Ac_{s}^{2}}(c_{s}^{2}+1)\right]-\frac{6}{r^{2}}e^{2\Lambda}-4\Psi^{\prime 2},$ (2.31)
where $c_{s}^{2}=\frac{dp_{r}}{d\rho}$ and $A=\frac{dp_{t}}{dp_{r}}$. Additional information can be found in Ref. [43] and the references therein.
For strange quark stars, where the surface energy density remains nonzero, it becomes necessary to introduce a correction to $y(R)$. Because of the discontinuity present in the energy distribution, the boundary condition is adjusted as follows: [48, 49, 50, 51]
$y_{R}\longrightarrow y_{R}-\frac{4\pi R^{3}\rho_{s}}{M},$ (2.32)
with $\rho_{s}$ being the difference in energy density across the stellar surface.
3 Equation of state and anisotropic profile
3.1 Equation of state
To investigate the impact of anisotropy on the fundamental nonradial oscillation modes, we chose to use the thermodynamically consistent vector MIT bag model, where the quark interaction is mediated by the vector channel $V^{\mu}$, as we can see in Ref. [52, 53]. Thus, the Lagrangian density is given by:
$\mathcal{L}_{\rm vMIT}=\left[\bar{\psi}_{q}\big{[}\gamma^{\mu}(i\partial_{\mu}-g_{qV}V_{\mu})-m_{q}\big{]}\psi_{q}-B+\frac{1}{2}m_{V}^{2}V^{\mu}V_{\mu}\right]\Theta(\bar{\psi}_{q}\psi_{q}),$ (3.1)
where $m_{q}$ is the mass of the quark (q)=(u,d,s)(q)=(u,\,d,\,s), $B$ denotes the pressure generated by the vacuum within the “bag” that confines the quarks, $\psi_{q}$ is the Dirac quark field, and $\Theta(\bar{\psi}_{q}\psi_{q})$ is the Heaviside step function. In the case where T=0[K]T=0\,[K], we can determine the total energy density:
$\rho=\sum_{q}\epsilon_{q}+B-\frac{1}{2}m_{v}^{2}V_{0}^{2},$ (3.2)
and the total pressure,
$p_{r}=\sum_{q}\mu_{q}n_{q}-\rho.$ (3.3)
The quantity $\epsilon_{q}$ is the energy density of the quarks:
$\epsilon_{q}=\frac{N_{c}}{\pi^{2}\hbar^{3}}\int_{0}^{k_{F_{q}}}\mu_{q}k^{2}d^{3}k,$ (3.4)
where $N_{c}=3$ is the number of colors, $k_{F_{q}}$ is the Fermi momentum of the quark $q$ and $\mu_{q}$ is the chemical potential. The parameters utilized in this work are the same as presented in Ref. [52]. We use $m_{u}=m_{d}=4\,[\rm MeV]$, and $m_{s}=95\,[\rm MeV]$. We also assume a universal coupling of quarks with the vector meson, i.e., $g_{uV}=g_{dV}=g_{sV}=g_{V}$, and use some values of $G_{V}$; as defined below
$G_{V}=\bigg{(}\frac{g_{V}}{m_{V}}\bigg{)}^{2},$ (3.5)
in units of $\rm fm]^{2}$. The value of the bag constant is taken as $B=81.1\,[\rm MeV/fm^{3}]$ and $G_{V}=0.1\,[\rm fm]^{2}$.
3.2 Anisotropic profile
To describe the anisotropic profile, we employ the quasilocal anisotropic profile. As described in [54], it depends on the fluid and spacetime variables of the form $\sigma=\sigma(p_{r},g_{11})$. Thus, this equation is given by:
$\sigma=\alpha p_{r}\left(1-\frac{1}{g_{11}}\right),$ (3.6)
where $\alpha$ is a dimensionless anisotropic constant and $g_{11}=e^{2\Lambda}$. The anisotropy model (3.6) was used, e.g., to analyze the effects of the anisotropy on the radial pulsation modes of polytropic stars [54, 55] and strange quark stars [56], nonradial oscillation modes within Cowling approximation of neutron stars [42] and strange quark stars [43], magnetic field structure [57], and slowly rotating neutron stars [58, 59]. To compare our results with some results reported in the literature -see, e.g., [42, 43]- we consider $2\leq\alpha\leq 2$.
For the anisotropic profile (3.6), we note that
$\frac{\partial\sigma}{\partial g_{11}}=\frac{\alpha p_{r}}{e^{4\Lambda}},\quad\frac{\partial\sigma}{\partial p_{r}}=\alpha\left(1-\frac{1}{g_{11}}\right),\quad{\rm and}\quad\frac{\partial\sigma}{\partial\rho}=\frac{\partial p_{r}}{\partial\rho}\frac{\partial\sigma}{\partial p_{r}}.$ (3.7)
It is important to highlight these factors since they appear in Eqs. (2.17)–(2.22). In this way, the set of equations to analyze nonradial oscillations is complete.
4 Anisotropic effects on the equilibrium, frequency of oscillations, and tidal deformability
4.1 General remarks
For nonradial oscillations mode: To determine the fundamental mode frequency, we choose a suitable trial Newtonian value $\omega_{Newt}=\sqrt{(M/R^{3})2l(l-1)/(2l+1)}$ as a starting point, as made in Ref. [40]. Then we start the integration process within the stellar interior. First, we construct three linearly independent solutions that satisfy the regularity conditions at the star’s center and numerically integrate them outward from $r=0$ to the midpoint at $R/2$. Next, we construct two additional linearly independent solutions that meet the boundary conditions at the stellar surface, and integrate them inward from $r=R$ to $R/2$. Finally at $R/2$, we combine these five solutions, in a way that ensures the resulting function satisfies the boundary conditions at both the center and the surface, completing the computation of the mode within the star.
The next step is to solve the Zerilli equation in the exterior region of the star. To accomplish this, we first determine the boundary values at the stellar surface for both the Zerilli function and its radial derivative. These can be extracted from the values of the metric perturbation functions $H(R)$ and $K(R)$ obtained through the interior integration. With these boundary conditions at $r=R$, the Zerilli equation is integrated outward in terms of the tortoise coordinate $r^{*}$. In the asymptotic regime, where $r^{*}\to\infty$, the general solution of the Zerilli equation can be represented as a linear combination of ingoing and outgoing waves, as can be found in Ref. [60].
To determine the quasi-normal mode frequencies, we have to integrate the Zerilli equation from the surface of the star to a sufficiently large radial coordinate, typically taken as $r_{\infty}\approx 50\,\omega^{-1}$ (During that integration, $\omega$ is treated as a real parameter in both the interior and exterior regions). The physical boundary condition at infinity imposes that only outgoing gravitational waves should be present. Therefore, the problem of finding quasi-normal mode frequencies reduces to locating the complex roots as done in [61].
This procedure is repeated iteratively: the real part of the estimated root is used as the new trial input in the next integration cycle. The iterations continue until the real part of the frequency converges to within one part in $10^{8}$ between successive steps. Additionally, once convergence is reached, the imaginary part of the frequency $\mathrm{Im}(\omega)$ yields the damping timescale of the mode.
For tidal deformability: To examine how anisotropy influences the tidal deformability of strange stars, we numerically integrate the stellar structure equations (2.5)-(2.7) together with the tidal deformability equation (2.29) from the center of the star ($r=0$) to its surface ($r=R$). Specifically, Eqs. (2.5)-(2.7) are first solved using the fourth-order Runge–Kutta method for various choices of the anisotropy parameter $\alpha$ and central energy density $\rho_{c}$. This step provides the radial profiles for $p_{r}$, $\sigma$, $\rho$, $m$, and $\Psi$. Next, Eq. (2.7) is handled via the shooting method: an initial trial for $\Psi_{c}$ at the center is supplied, and if the resulting solution does not satisfy the boundary condition given in equation (2.7), $\Psi_{c}$ is iteratively adjusted until convergence is achieved. With the consistent $\Psi_{c}$ determined, the tidal deformability equation (2.29) is then integrated outward from $r=0$ to $r=R$, ensuring a self-consistent solution for each selected pair of $\alpha$ and $\rho_{c}$. Once the deformability function $y(r)$ is known, the value of $y(r=R)$ is determined by means of Eq. (2.3), which finally allows us to calculate the dimensionless deformability parameter (2.27).
4.2 Equilibrium configurations and frequency of oscillations
Figure 1: The total gravitational mass, in Sun’s masses, against the central energy density and as a function of the total radius are plotted on the left and right panels, respectively. In both panels, five different anisotropic parameters $\alpha$ are employed. The solid purple and black curves indicate the $95\%$ confidence intervals for the masses and radii of PSR J0030+04510030+0451 [62, 63], while the orange and gray curves show the same for PSR J0740+66200740+6620 [64, 65], both measured by the NICER collaboration. In all cases, the outer lines offer a greater likelihood of containing the true value but with lower precision, whereas the inner lines provide higher accuracy but with lower statistical coverage.
The normalized mass with Solar mass as a function of the central energy density and versus the radius is presented on the left and right panels of Fig. 1, respectively, for five different anisotropic parameters. On the right panel, the masses-radii relation reported by NICER from compact stars PSR J0030+04510030+0451 [62, 63] and PSR J0740+66200740+6620 [64, 65] are presented. The corresponding bands of the pulsars PSR J0740+66200740+6620 [66], PSR J0348+04320348+0432 [67], and PSR J1614+22301614+2230 [68], respectively marked by light yellow, light orange, and light green colors, are also shown in the mass-radius diagram. In such figure, since the central pressure becomes extremely high or low for large and small values of the anisotropic parameter, respectively, it is impossible to carry out numerical calculations with high central energy densities. Thus, we consider all equilibrium configurations where excellent numerical precision is found. Therefore, in such situations, the approach is to get as close as possible to the value corresponding to the maximum mass. In all mass against central energy density curves, the mass increases monotonically with $\rho_{c}$ until values near the maximum mass values. In all mass versus radius curves, the mass grows with the radius proportioning to $M\propto R^{3}$ until it reaches the maximum radius. At this point, all curves turn anti-clockwise, so that the masses start to increase with the diminution of the radius until it attains the maximum-mass value.
Figure 2: The $f$-mode frequency of oscillations and the normalized $\omega_{f}$ with the average density $(M/R^{3})^{0.5}$ against the total mass are presented on the left and right panels, respectively. In both panels, five different anisotropic parameters $\alpha$ are used.
The effect of anisotropic pressure on mass and radius can also be observed in Fig. 1. When examining the maximum mass and its corresponding total radius, it becomes evident that these quantities vary significantly with changes in $\alpha$. Once $\alpha$ exceeds zero, the growth of $\alpha$ requires the star to attain a higher maximum mass and a larger radius to achieve equilibrium. Moreover, the increase in mass and radius with $\sigma$ can be understood by noting the change of $\sigma$ with the parameter $\alpha$. Greater anisotropy $\sigma$ supports higher fluid pressure, enabling stars with larger masses and radii to resist collapse. Conversely, lower anisotropy leads to equilibrium states characterized by smaller mass and radius.
The $f$-mode nonradial oscillation frequency and the eigenfrequency normalized by the square root of the average density, $(\sqrt{M/R^{3}})$, as functions of the total mass, are shown in the left and right panels of Fig. 2, respectively, for selected values of the anisotropy parameter $\alpha$.
In the left panel, all curves show that the $f$-mode frequency increases monotonically with the total mass up to the maximum-mass configuration. The influence of anisotropy on the fluid pulsation mode is also evident. From the $f$-mode frequency versus $M/M_{\odot}$ curves, in certain mass ranges, the $f$-mode frequency increases or decreases in correspondence with an increase or decrease in $\alpha$.
In the right panel, for the isotropic case ($\alpha=0$), the relationship between $\omega_{f}\sqrt{R^{3}/M}$ and $M/M_{\odot}$ is nearly linear. When $\alpha<0$, the normalized eigenfrequency $f$ decreases monotonically as $M/M_{\odot}$ increases. For $\alpha>0$, in some mass ranges, the normalized oscillation eigenfrequency $f$ decreases with increasing total mass until it reaches a minimum value of $\omega_{f}\sqrt{R^{3}/M}$. Beyond this point, the curve bends counterclockwise, and $\omega_{f}\sqrt{R^{3}/M}$ starts to increase with $M/M_{\odot}$ until the maximum-mass configuration is reached. As noted in [56], variations in the oscillation frequency are linked to changes in the radial fluid pressure induced by anisotropy.
$\rho_{c}\,[\rm MeV/fm^{3}]$ $\alpha$ $M/M_{\odot}$ $R\,[\rm km]$ $M/R$ $f_{f}\,[\rm kHz]$ $\omega_{f}\,\sqrt{R^{3}/M}$
1.0-1.0 $0.5567$ $7.5531$ $0.1086$ $1.9301$ $0.9321$
0.5-0.5 $0.5950$ $7.7154$ $0.1139$ $1.9290$ $0.9279$
$400$ $0.0$ $0.6388$ $7.8925$ $0.1195$ $1.9279$ $0.9237$
+0.5+0.5 $0.6913$ $8.0870$ $0.1262$ $1.9262$ $0.9229$
+1.0+1.0 $0.7504$ $8.3013$ $0.1335$ $1.9242$ $0.9211$
1.0-1.0 $1.1918$ $9.2840$ $0.1895$ $2.1073$ $0.9437$
0.5-0.5 $1.3055$ $9.5340$ $0.2022$ $2.1032$ $0.9317$
$600$ $0.0$ $1.4386$ $9.8033$ $0.2167$ $2.0976$ $0.9252$
+0.5+0.5 $1.5964$ $10.0913$ $0.2336$ $2.0895$ $0.9135$
+1.0+1.0 $1.7802$ $10.3940$ $0.2528$ $2.0781$ $0.9020$
Table 1: The mass $M/M_{\odot}$, radius $R$, compactness $M/R$, $f$-mode oscillation frequency, and eigenfrequency normalized with the mean density $\omega_{f}\,\sqrt{R^{3}/M}$ for strange quark stars with two different central energy densities and different values of $\alpha$.
Table 1 presents the mass, radius, compactness, $f$-mode pulsation frequency, and the oscillation eigenfrequency normalized with the mean density, for compact objects for the central energy densities $400\,[\rm MeV/fm^{3}]$ and $600\,[\rm MeV/fm^{3}]$ and five values of anisotropic parameters $\alpha$. It can be noted that, when varying $\alpha$ from 1.0-1.0 to +1.0+1.0, the mass, radius, compactness, $f$-mode, and the normalized oscillation eigenfrequency change respectively in almost $35\%$, $10\%$, $\%, 3.2-3.2\%$, and $1.1\%$ for $\rho_{c}=400\,[\rm MeV/fm^{3}]$ and in $49\%$, $12\%$, $33\%$, $1.4\%$, and $4.4\%$ for $\rho_{c}=600\,[\rm MeV/fm^{3}]$. Thus, for a central energy density interval, we note that when $\alpha$ increases, the mass, radius, and compactness rise, whereas the $f$-mode frequency (and the normalized eigenfrequency $\omega_{f}$) decreases. From these results, together with those shown in Figs. 1 and 2, we see that as the star’s mass and radius increase, the oscillation mode $f$ decreases. This peculiarity has also been found in the study of compact stars in other contexts. For example, in studies of dark matter admixed with hadronic matter [69, 20] and quark matter [70, 16], in hybrid stars with hyperons and delta baryons [71], and in compact stars with crust [14].
4.3 On the detectability of the fundamental mode signal
A core-collapse supernova (CCSN) is a violent explosion marking the end of the life of a massive star with $M\gtrsim 8M_{\odot}$. Once the stellar core exceeds the effective Chandrasekhar limit ($\sim 1.4M_{\odot}$), it becomes gravitationally unstable and collapses to nuclear densities. The ensuing collapse drives a shock wave that triggers the CCSN explosion. Gravitational waves and neutrinos are emitted directly from the collapsing core, providing unique probes of the physical mechanisms underlying CCSN dynamics. Following the explosion, a hot proto-neutron star may form; during its early evolution, the gravitational-wave spectrum is dominated by the fundamental $f$-mode and the first pressure mode $p_{1}$ [72]. A typical CCSN releases a total energy of $\sim 10^{53}$ erg ($\sim 0.056M_{\odot}$), and part of this energy may be radiated through $f$-mode oscillations [72, 73].
Let us consider the detection of gravitational waves associated with the fundamental oscillation mode of a magnetar formed after a supernova explosion. It is well established that when a gravitational wave reaches the detector, the signal takes the following form:
$h(t)=\textit{h}e^{-t/\tau}{\rm sin}[2\pi ft],$ (4.1)
with f being the fundamental mode frequency, $\tau$ is the damping time and the gravitational wave amplitude $h$ is given by
$\textit{h}\sim A\left(\frac{E_{gw}}{10^{-6}M_{\odot}}\right)^{1/2}\left(\frac{10{\rm kpc}}{d}\right)\left(\frac{1{\rm kHz}}{f}\right)\left(\frac{{1\rm ms}}{\tau}\right)^{1/2},$ (4.2)
where $A=2.4\times 10^{-20}$, $E_{gw}$ is the energy released through the fundamental mode and $d$ is the distance to the source [74, 73].
The signal-to-noise ratio at the detector reads [74, 73]
$\left(\frac{S}{N}\right)^{2}=\frac{4Q_{\textit{F}}^{2}}{1+4Q_{\textit{F}}^{2}}\;\frac{h^{2}\tau}{2S_{n}}.$ (4.3)
Here, $Q_{\textit{F}}\equiv\pi f\tau$ is the quality factor and $S_{n}$ is the noise power spectral density.
From Eqs. (4.2) and (4.3), the energy radiated by the fundamental mode is given by
$\left(\frac{E_{\rm gw}}{M_{\odot}}\right)=BC\left(\frac{S}{N}\right)^{2}\left(\frac{d}{10{\rm kpc}}\right)^{2}\left(\frac{f}{1{\rm kHz}}\right)^{2}\left(\frac{S_{n}}{1{\rm Hz^{-1}}}\right),$ (4.4)
where $B=3.47\times 10^{36}$ and $C=(1+4Q_{\textit{F}}^{2})/4Q_{\textit{F}}^{2}$. In the following, we estimate the minimum energy that must be emitted in order to achieve a signal-to-noise ratio (S/N) greater than $5$, for a source that could be observed within our galaxy, i.e, at $d\sim 10\,[\rm kpc]$
$\rho_{c}\,[\rm MeV/fm^{3}]$ $\alpha$ $\tau\,[\rm ms]$ $E^{LV}_{gw}\,[\rm 10^{-7}M_{\odot}]$ $E^{ET}_{gw}\,[\rm 10^{-10}M_{\odot}]$
1.0-1.0 $577.477$ $1.2926$ $3.2316$
0.5-0.5 $530.746$ $1.2912$ $3.2280$
$400$ $0.0$ $485.668$ $1.2897$ $3.2243$
+0.5+0.5 $443.167$ $1.2874$ $3.2186$
+1.0+1.0 $402.307$ $1.2847$ $3.2119$
1.0-1.0 $187.033$ $1.5409$ $3.8523$
0.5-0.5 $174.550$ $1.5349$ $3.8373$
$600$ $0.0$ $164.209$ $1.5267$ $3.8169$
+0.5+0.5 $157.036$ $1.5150$ $3.7875$
+1.0+1.0 $153.927$ $1.4985$ $3.7462$
Table 2: The damping time $\tau$ for strange quark stars, the energy required to excite the fundamental mode, $E^{LV}_{gw}$ and $E^{ET}_{gw}$, for the signal to be detectable by Advanced LIGO-Virgo and the Einstein Telescope, respectively. For this result, we considered the source at $d\sim 10$ $\rm[\rm kpc]$, a signal-to-noise ratio S/N $\sim 5$, two different central densities, and different values of $\alpha$.
In Table 2, we show the minimum energy required for the fundamental mode to be detected. For this objective, we consider two detectors; the first one is Advanced LIGO-Virgo with a sensitivity of $S_{n}^{1/2}\sim 2\times 10^{-23}\,\mathrm{[\rm Hz]}^{-1/2}$ at $\sim[\rm kHz]$ [1], and the second one is the Einstein telescope with a sensitivity $S_{n}^{1/2}\sim 10^{-24}\,\mathrm{[\rm Hz]}^{-1/2}$ at similar frequency band [75].
Our results show that for stars of about $1.4\,M_{\odot}$ the $f$-mode frequency is near to $2.0$ kHz and the damping time is about $164.0$ ms. Also, as can be seen, at a distance of $10$ kpc, the minimum energy emitted in the GW must be $1.53\times 10^{-7}M_{\odot}$. As can be seen, this required energy is much lower than the energy released in a CCSN explosion, which is about ($10^{-5}-10^{-6}M_{\odot}$ ).
For a fixed central energy density $\rho_{c}=600\,[\rm MeV/fm^{3}]$, varying the anisotropy from $\alpha=-1.0$ to $\alpha=+1.0$ produces measurable shifts in the $f$-mode properties and the GW energy needed for detectability. For $\alpha=-1.0$ we obtain $M=1.1918\,M_{\odot}$, $f=2.1073\,[\rm kHz]$, $\tau=187.033\,[\rm ms]$, with detection thresholds $E^{LV}_{\mathrm{gw}}>1.5409\times 10^{-7}M_{\odot}$ (Advanced LIGO/Virgo, S/N>5>5) and $E^{ET}_{\mathrm{gw}}>3.8523\times 10^{-10}M_{\odot}$ (Einstein Telescope, S/N>5>5). For $\alpha=+1.0$ we find $M=1.7802\,M_{\odot}$, $f=2.0781\,[\rm kHz]$, $\tau=153.927\,[\rm ms]$, and thresholds $E^{LV}_{\mathrm{gw}}>1.4985\times 10^{-7}M_{\odot}$ and $E^{ET}_{\mathrm{gw}}>3.7462\times 10^{-10}M_{\odot}$. This systematic response of $(f,\tau,E_{\mathrm{gw}})$ to $\alpha$-combined with independent constraints on $M$ and $R$-provides a practical diagnostic for future GW asteroseismology: precise measurements can bound the level of anisotropy and, given the higher compactness and typically higher $f$-mode frequencies expected for strange-matter stars, help discriminate hadronic from strange-matter compositions.
4.4 Tidal deformability
Figure 3: left: Dimensionless tidal deformability against the total mass for different values of $\alpha$. right: Oscillation frequency $f_{f}$ as a function of the dimensionless tidal deformability for five values of $\alpha$. In both panels, the vertical dashed line depicts dimensionless tidal deformability $\Lambda_{1.4}=190^{+390}_{-120}$ from the event GW170817170817 reported by LVC in Ref. [76].
Fig. 3 plots, in its left and right panels respectively, the dimensionless tidal deformability $\Lambda$ against the total mass $M/M_{\odot}$ and the oscillation frequencies of the fundamental mode $f_{f}$ as a function of the dimensionless tidal deformability $\Lambda$ for several choices of the anisotropy parameter $\alpha$. These curves are compared with the observational range for $\Lambda_{1.4}=190^{+390}_{-120}$ reported by LIGO-Virgo [76]. On the left panel, in all cases, the tidal deformability declines steadily as the total mass increases. Additionally, the influence of anisotropy on the deformability response is clearly seen: when $\alpha$ is positive, the calculated $\Lambda$ values are larger for a given mass, whereas negative $\alpha$ leads to smaller values. Importantly, all curves remain within the $\Lambda_{1.4}$ band determined by LIGO-Virgo. On the right panel, the plots show that, as the tidal deformability increases, the $f$-mode frequency decreases monotonically.
Figure 4: The dimensionless tidal deformabilities of the GW170817170817 event components are shown for different values of the anisotropic parameter $\alpha$. The yellow line depicts the LIGO-Virgo confidence curves [1], while the dotted diagonal line marks the points where $\Lambda_{1}=\Lambda_{2}$.
Using the observational data provided by LVC, the study in [1] established bounds on $\Lambda_{1}$ and $\Lambda_{2}$, which quantify the dimensionless tidal deformability of the binary system. Here, $\Lambda_{1}$ corresponds to the tidal deformability of the more massive star, while $\Lambda_{2}$ refers to that of its companion. Fig. 4 presents the $\Lambda_{1}$$\Lambda_{2}$ plane, where each curve is generated by selecting a specific value of $M_{1}$ and then computing $M_{2}$ using the chirp mass $\mathcal{M}=1.188\,M_{\odot}$ [1], defined as
$\mathcal{M}=\frac{(M_{1}M_{2})^{3/5}}{(M_{1}+M_{2})^{1/5}}.$ (4.5)
The stellar masses considered fall within the intervals $1.36\leq M_{1}/M_{\odot}\leq 1.60$ and $1.17\leq M_{2}/M_{\odot}\leq 1.36$, respectively. In addition, the $50\%$ and $90\%$ credibility contours for the GW170817170817 event, provided by LVC under the low-spin assumption, are included in the plot. For both $\alpha>0$ and $\alpha<0$, the effect of the anisotropy parameter on the tidal deformability becomes evident. Furthermore, we can observe that larger positive values of $\alpha$ help to have dimensionless tidal deformability values closer to the confidence intervals reported in [1].
5 Conclusions
This article reports the effects of anisotropy on both the stellar structure and $f$ oscillation modes of strange stars. For this aim, we used the stellar equilibrium equations, including the anisotropic factor, and we derived the nonradial oscillation equations involving such a factor within the fully general relativity context. For the fluid contained in the star, we assume that it follows the MIT bag model with vector coupling, and the anisotropy is described by the profile $\sigma=\alpha p_{r}(1-\frac{1}{g_{11}})$.
Regarding the equilibrium configurations, we found that the strange star is affected by anisotropy. For certain ranges of central energy density, we observed both higher (lower) mass and radius values when using higher (lower) values of $\alpha$; this is consistent with what has been reported in the literature for strange stars using the MIT bag model EOS. For $\alpha>0$, increasing $\alpha$ favors the proposed model, bringing the masses closer to the observational data. This suggests that anisotropy could play a relevant role in identifying strange stars.
In our study of nonradial oscillation modes, we found that the $f$-modes are strongly influenced by anisotropy. For specific ranges of total mass, we observed that the $f$-modes increase (or decrease) as the anisotropy parameter $\alpha$ increases (or decreases). It is worth noting that the equations governing nonradial oscillations depend on the chosen anisotropic profile; in this case, they are determined by the form of equation (3.6).
We also investigated the impact of anisotropy on the detectability of the fundamental oscillation mode. Our results indicate that massive stars with an anisotropy parameter $\alpha\sim 1.0$ could be detected within our galaxy, i.e., at distances of approximately $\sim 10\,[\rm kpc]$.
Finally, we also examine the consistency between the dimensionless tidal deformability of anisotropic strange stars and the observational constraints reported by the LVC for the GW170817170817 event. Our analysis shows that the results presented in this work fall within the observational bounds provided by LVC. Our findings indicate that the tidal deformability increases for positive values of $\alpha$ and decreases for negative values (-$\alpha$). Moreover, several studies in the literature have explored the tidal deformability of strange stars under various theoretical frameworks. For instance, [51] investigates this parameter considering quark matter in the color-flavor-locked (CFL) phase of color superconductivity, [77] examines the role of isospin effects in strange quark matter, and [78] analyzes the tidal deformability assuming a quasiparticle model that incorporates nonperturbative aspects of quantum chromodynamics in the low-density regime. In all these works, as in the present study, constraints from the GW170817170817 event are used to place limits on the properties of strange stars within the corresponding theoretical scenarios.
Acknowledgments
JDVA thanks Universidad Privada del Norte and Universidad Nacional Mayor de San Marcos for the financial support - RR No. 005753\,005753-$2021$-R//UNMSM under the project number B2113178121131781. JMZP acknowledges support from FAPERJ, Process SEI-260003/000308/2024. CVF acknowledges the financial support of the productivity program of the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), with Project No. 304569/2022-4. CHL is thankful to the São Paulo Research Foundation FAPESP (Grant No. 2020/05238-9) and to CNPq (Grants No.401565/2023-8 and 305327/2023-2).
Appendix A Nonradial perturbations of relativistic spherically symmetric stars in the presence of an anisotropic fluid
A.1 Perturbative variables
For even-parity harmonics, the equations of motion are found through the perturbed Einstein’s field equation (2.12), by employing the metric perturbation (2.16) and the Lagrangian fluid displacement vector components in the form
$\xi^{r}=\frac{1}{r^{2}}e^{-\Lambda}WY^{\ell m},$ (A.1)
$\xi^{\theta}=-\frac{1}{r^{2}}V\partial_{\theta}Y_{\ell m},$ (A.2)
$\xi^{\phi}=-\frac{1}{r^{2}\sin^{2}\theta}V\partial_{\phi}Y_{\ell m},$ (A.3)
where $W$ and $V$ are functions be dependent on $t$ and $r$. Moreover, the non-perturbed four-velocity and four-radial unit vectors are used, which are respectively placed as
$u^{(0)\mu}=\left(\frac{1}{e^{\Psi}},0,0,0\right)\quad{\rm and}\quad k^{(0)\mu}=\left(0,\frac{1}{e^{\Lambda}},0,0\right),$ (A.4)
respectively. The perturbed four-velocity components are set in the form
$\delta u^{t}=\frac{1}{2}\frac{HY_{\ell m}}{e^{\Psi}},$ (A.5)
$\delta u^{r}=\frac{1}{e^{\Psi}}\frac{d\xi^{r}}{dr}=\frac{\dot{W}Y_{\ell m}}{r^{2}e^{\Psi+\Lambda}},$ (A.6)
$\delta u^{\theta}=\frac{1}{e^{\Psi}}\frac{d\xi^{\theta}}{dr}=-\frac{\dot{V}\partial_{\theta}Y_{\ell m}}{r^{2}e^{\Psi}},$ (A.7)
$\delta u^{\phi}=\frac{1}{e^{\Psi}}\frac{d\xi^{\phi}}{dr}=-\frac{\dot{V}\partial_{\phi}Y_{\ell m}}{r^{2}e^{\Psi}\sin^{2}\theta},$ (A.8)
and the perturbed radial unit vectors given by
$\delta k^{0}=\frac{1}{e^{\Psi}}\left[\frac{H_{1}}{e^{\Psi+\Lambda}}+\frac{\dot{W}}{r^{2}e^{\Psi}}\right]Y_{\ell m}\quad{\rm and}\quad\delta k^{1}=-\frac{1}{2}\frac{H}{e^{\Lambda}}Y_{\ell m}.$ (A.9)
The relations defined in Eqs. (A.4), (A.5), and (A.9) satisfy the equalities $g_{\mu\nu}k^{\mu}k^{\nu}=1$, $g_{\mu\nu}u^{\mu}k^{\nu}=0$, and $g_{\mu\nu}u^{\mu}u^{\nu}=0$ in their both non-perturbed and perturbed form.
A.2 Perturbation equations
To obtain the equations of motion of the pulsating configuration, following Ref. [38], the Lagrangian perturbation of the baryon number density $N$ is
$\frac{\Delta N}{N}=\xi^{k}_{\,\,|k}-\frac{1}{2}\frac{\delta\left[{}^{(3)}g\right]}{{}^{(3)}g}$. (A.10)
The term $\xi^{k}_{\,\,|k}$ denotes the covariant derivative in the $3$-dimensional geometry at constant time, and the terms $\delta\left[{}^{(3)}g\right$] and $^{(3)}g$ depict the determinant of the perturbed and unperturbed $3$-dimensional metric, respectively. By employing the background metric (2.4), the metric perturbation (2.16), and the fluid displacement vectors (A.1)-(A.3), the equation (A.10) yields
$\frac{\Delta N}{N}=-\frac{1}{2}HY_{\ell m}-\frac{e^{-\Lambda}}{r^{2}}W^{\prime}Y_{\ell m}-\frac{V}{r^{2}}\ell(\ell+1)Y_{\ell m}-KY_{\ell m}.$ (A.11)
Using the baryon conservation equation, $\nabla_{\mu}(Nu^{\mu})=0$, the local law of energy conservation ($u^{\nu}\nabla^{\mu}T_{\mu\nu}$=0) in terms of Lagrangian perturbations yield:
$\Delta\rho=(\rho+p_{r}+\sigma)\frac{\Delta N}{N}.$ (A.12)
Employing the relations of the thermodynamic functions in their Lagrangian and Eulerian forms, $\Delta\rho\thickapprox\delta\rho+\frac{\partial\rho^{0}}{\partial r}\xi^{r}$, and equation (A.11), we find that equation (A.12) takes the form:
$\hskip-28.45274pt\delta\rho=-(p_{r}+\rho+\sigma)\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]Y_{\ell m}-\rho^{\prime}\frac{e^{-\Lambda}}{r^{2}}WY_{\ell m}.$ (A.13)
The perturbed radial pressure can be determined using the EOS $p_{r}=p_{r}(\rho)$, where $\delta p_{r}=\frac{dp_{r}}{d\rho}\delta\rho$. From Eq. (2.12), the nonnull perturbed Einstein field equations are given by
$H^{\prime}+\frac{He^{2\Lambda}}{r}\left[4\pi r^{2}(p_{r}-\rho+\sigma)+1+\frac{\ell(\ell+1)}{2}\right]=rK^{\prime\prime}-\frac{8\pi\rho^{\prime}e^{\Lambda}W}{r}-\frac{8\pi e^{2\Lambda}V\ell(\ell+1)}{r}\left(p_{r}\right$.
$\left.+\rho+\sigma\right)-\frac{8\pi e^{\Lambda}W^{\prime}}{r}\left(p_{r}+\rho+\sigma\right)-\frac{Ke^{2\Lambda}}{r}\left[8\pi r^{2}(p_{r}+\rho+\sigma)-1+\frac{\ell(\ell+1)}{2}\right]-e^{2\Lambda}\left[4\pi\rho r^{2}\right$.
$\left.+\frac{5m}{r}-3\right]K^{\prime},$ (A.14)
$\hskip-11.38092pt\ell(\ell+1)H_{1}=2r^{2}{\dot{K}}^{\prime}+2r{\dot{K}}e^{2\Lambda}\left(-4\pi r^{2}p_{r}-\frac{3m}{r}+1\right)-2r{\dot{H}}-16\pi(p_{r}+\rho)e^{\Lambda}{\dot{W}},$ (A.15)
$\hskip-11.38092ptH_{1}^{\prime}=-\frac{2e^{\Lambda}H_{1}}{r}\left[\frac{m}{r}+2\pi r^{2}(p_{r}-\rho)\right]+e^{2\Lambda}{\dot{H}}+e^{2\Lambda}{\dot{K}}+16\pi\left(\rho+p_{r}+\sigma\right)e^{2\Lambda}{\dot{V}},$ (A.16)
$\hskip-11.38092pt{\ddot{K}}-e^{2\Psi-2\Lambda}K^{\prime\prime}-\frac{Ke^{2\Psi}}{r^{2}}\left(8\pi r^{2}\left(\Gamma p_{r}-\rho-p_{r}-\sigma+\sigma\frac{dp_{r}}{d\rho}\right)-\ell(\ell+1)+2\right)$
$\hskip-11.38092pt-\frac{2K^{\prime}e^{2\Psi}}{r}\left(1-\frac{m}{r}+2\pi r^{2}(p_{r}-\rho)\right)-\frac{e^{2\Psi}H}{r^{2}}\left(4\pi(p_{r}+\rho-\sigma+\Gamma p_{r})r^{2}-2+\frac{4m}{r}+4\pi\sigma r^{2}\frac{dp_{r}}{d\rho}\right)$
$\hskip-11.38092pt+\frac{W^{\prime}e^{2\Psi-\Lambda}}{r^{2}}\left(8\pi\left(\rho+p_{r}+\sigma-\Gamma p_{r}\right)-8\pi\sigma\frac{dp_{r}}{d\rho}\right)+\frac{Ve^{2\Psi}}{r^{2}}8\pi\ell(\ell+1)\left(\rho+p_{r}+\sigma-\Gamma p_{r}-\sigma\frac{dp_{r}}{d\rho}\right)$
$+W\left(\rho^{\prime}-p_{r}^{\prime}\right)\frac{8\pi e^{2\Psi-\Lambda}}{r^{2}}=0,$ (A.17)
${\dot{H}}_{1}=H^{\prime}e^{2\Psi}-K^{\prime}e^{2\Psi}+\frac{He^{2\Psi+2\Lambda}}{r}\left(\frac{2m}{r}+8\pi r^{2}p_{r}\right).$ (A.18)
From the perturbed conservation law of the energy-momentum tensor components $\delta\left(\nabla_{\nu}T^{\nu}_{1}\right)$ and $\delta\left(\nabla_{\nu}T^{\nu}_{2}\right)$, we respectively obtain:
$\hskip-11.38092pt-\frac{H^{\prime}}{2}(p_{r}+\rho)e^{\Psi}-e^{\Psi}\Psi^{\prime}(p_{r}+\rho)\left[\frac{H}{2}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]-e^{\Psi}\Psi^{\prime}\sigma\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}\right$.
$\hskip-11.38092pt\left.+\frac{V}{r^{2}}\ell(\ell+1)+K\right]-\left\{e^{\Psi}\left(p_{r}+\rho+\sigma\right)\frac{dp_{r}}{d\rho}\left[\frac{H}{2}+\frac{e^{-\Lambda}}{r^{2}}W^{\prime}+\frac{V}{r^{2}}\ell(\ell+1)+K\right]\right\}^{\prime}-K^{\prime}\sigma e^{\Psi}$
$\hskip-11.38092pt+(p_{r}+\rho)\frac{e^{\Lambda-\Psi}}{r^{2}}{\ddot{W}}+(p_{r}+\rho)e^{\Psi}\left[H^{\prime}-K^{\prime}+\frac{H}{r}e^{2\Lambda}\left(\frac{2m}{r}+8\pi r^{2}p_{r}\right)\right]+(p_{r}+\rho)\left(\Psi^{\prime 2}+\Psi^{\prime\prime}\right$.
$\hskip-11.38092pt\left.-\Psi^{\prime}\Lambda^{\prime}\right)\frac{e^{\Psi-\Lambda}}{r^{2}}W-2(p_{r}+\rho+\sigma)\Psi^{\prime}\frac{e^{\Psi-\Lambda}}{r^{3}}W+\frac{2}{r^{3}}\left(\sigma\Lambda^{\prime}+\frac{3\sigma}{r}-\sigma^{\prime}\right)e^{\Psi-\Lambda}W-\frac{2\sigma}{r^{3}}e^{\Psi-\Lambda}W^{\prime}$
$+p_{r}^{\prime}\Psi^{\prime}\frac{e^{\Psi-\Lambda}}{r^{2}}W-\frac{2e^{\Psi}}{r}(\delta{\tilde{\sigma}})=0,$ (A.19)
$\hskip-11.38092pt(p_{r}+\rho+\sigma)\frac{\ddot{V}}{e^{2\Psi}}+\frac{H}{2}\left(p_{r}+\rho+\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+\frac{W^{\prime}}{r^{2}e^{\Lambda}}\left(\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+\frac{V\ell(\ell+1)}{r^{2}}\left(\Gamma p_{r}\right$.
$\hskip-11.38092pt\left.+\sigma\frac{dp_{r}}{d\rho}\right)+K\left(\Gamma p_{r}+\sigma\frac{dp_{r}}{d\rho}\right)+p_{r}^{\prime}\frac{W}{r^{2}e^{\Lambda}}-\delta{\tilde{\sigma}}=0,$ (A.20)
where $\delta\sigma=\delta{\tilde{\sigma}}Y_{\ell m}$. It is important to mention that the term $\delta{\tilde{\sigma}}$ is renamed as $\delta{\sigma}$. In addition, for the anisotropic profile $\sigma=\sigma(p_{r},g_{11})$, we found that
$\delta\sigma=\frac{\partial\sigma}{\partial p_{r}}\delta p_{r}+\frac{\partial\sigma}{\partial g_{11}}h_{11}.$ (A.21)
This last relation is also used to determine Eqs. (2.17)-(2.22).
Appendix B Perturbation functions near $r=0$
For the isotropic case, $\sigma=0$, the spacetime perturbation and fluid functions expanded in a power series near $r=0$ have been previously reported in Ref. [40]. This appendix shows these expanded functions for the anisotropic case, $\sigma\neq 0$. Since we seek solutions to the perturbation equations (2.18)–(2.21), the perturbation functions ${\tilde{H}}_{1}(r)$, ${\tilde{K}}(r)$, ${\tilde{W}}(r)$, and ${\tilde{X}}(r)$ near the center of the star are considered into the form
${\tilde{H}}_{1}(r)={\tilde{H}}_{1}(0)+\frac{1}{2}{\tilde{H}}_{1}^{\prime\prime}(0)r^{2}+O(r^{4}),$ (B.1)
${\tilde{K}}(r)={\tilde{K}}(0)+\frac{1}{2}{\tilde{K}}^{\prime\prime}(0)r^{2}+O(r^{4}),$ (B.2)
${\tilde{W}}(r)={\tilde{W}}(0)+\frac{1}{2}{\tilde{W}}^{\prime\prime}(0)r^{2}+O(r^{4}),$ (B.3)
${\tilde{X}}(r)={\tilde{X}}(0)+\frac{1}{2}{\tilde{X}}^{\prime\prime}(0)r^{2}+O(r^{4}),$ (B.4)
where the first-order -${\tilde{H}}_{1}(0)$, ${\tilde{K}}(0)$, ${\tilde{W}}(0)$, and ${\tilde{X}}(0)$- and second-order coefficients -${\tilde{H}}_{1}^{\prime\prime}(0)$, ${\tilde{K}}^{\prime\prime}(0)$, ${\tilde{W}}^{\prime\prime}(0)$, and ${\tilde{X}}^{\prime\prime}(0)$- are constants. The first-order terms are given by:
$\hskip-11.38092pt{\tilde{H}}(0)={\tilde{K}}(0),$ (B.5)
$\hskip-11.38092pt{\tilde{X}}(0)=(\rho_{{}_{0}}+p_{r{{}_{0}}})e^{\Psi_{0}}\left[-\frac{\omega^{2}e^{-2\Psi_{{}_{0}}}}{\ell}{\tilde{W}}(0)+\frac{4\pi}{3}\left(\rho_{{}_{0}}+3p_{r_{0}}\right){\tilde{W}}(0)-\frac{{\tilde{K}}(0)}{2}\right]+\alpha p_{r_{0}}\left[{\tilde{K}}(0)\right$.
$\left.-\frac{16\pi}{3}\rho_{{}_{0}}{\tilde{W}}(0)\right]e^{\Psi_{0}},$ (B.6)
$\hskip-11.38092pt{\tilde{H}}_{1}{(0)}=\left[\frac{2\ell{\tilde{K}}(0)-16\pi(\rho_{0}+p_{r_{0}}){\tilde{W}}(0)}{\ell(\ell+1)}\right],$ (B.7)
$\hskip-11.38092pt{\tilde{V}}(0)=-\frac{1}{\ell}{\tilde{W}}(0),$ (B.8)
with the constants $\rho_{{}_{0}}$, $p_{r}{{}_{0}}$, and $\Psi_{{}_{0}}$ representing the first-order terms of the power-series expansions:
$\rho(r)=\rho_{{}_{0}}+\frac{1}{2}\rho_{{}_{2}}r^{2},$ (B.9)
$p_{r}(r)=p_{r_{0}}+\frac{1}{2}p_{r_{2}}r^{2}+\frac{1}{4}p_{r_{4}}r^{4},$ (B.10)
$\Psi(r)=\Psi_{{}_{0}}+\frac{1}{2}\Psi_{{}_{2}}r^{2}+\frac{1}{4}\Psi_{{}_{4}}r^{4},$ (B.11)
$\sigma(r)=\sigma_{{}_{0}}+\frac{1}{2}\sigma_{{}_{2}}r^{2}+\frac{1}{4}\sigma_{{}_{4}}r^{4},$ (B.12)
and $\alpha$ being the dimensionless anisotropic constant. Through equation (3.6), we can see that in the star’s center, because $g_{11}=1$ (see Eq. (2.8)), we have $\sigma_{{}_{0}}=0$. The second and fourth-order coefficients of the power-series (B.9)-(B.12) are
$\hskip-11.38092pt\rho_{{}_{2}}=-\frac{(\rho_{{}_{0}}+p_{r_{0}})^{2}}{2\Gamma\,p_{r_{0}}}\Psi_{{}_{2}},$ (B.13)
$\hskip-11.38092ptp_{r_{2}}=-\frac{4\pi}{3}(\rho_{{}_{0}}+3p_{r_{0}})(\rho_{{}_{0}}+p_{r_{0}})+\frac{16\pi}{3}\alpha\rho_{{}_{0}}\,p_{r_{0}},$ (B.14)
$\hskip-11.38092pt\Psi_{{}_{2}}=\frac{8\pi}{3}(\rho_{{}_{0}}+3p_{r_{0}})-\frac{16\pi\alpha}{3}\frac{\rho_{{}_{0}}p_{r_{0}}}{\rho_{{}_{0}}+p_{r_{0}}},$ (B.15)
$\hskip-11.38092pt\sigma_{{}_{2}}=\frac{16\pi\alpha}{3}\rho_{{}_{0}}\,p_{r_{0}},$ (B.16)
$\hskip-11.38092ptp_{r_{4}}=-\frac{2\pi}{5}(\rho_{{}_{2}}+5p_{r_{2}})(\rho_{0}+p_{r_{0}})-\frac{2\pi}{3}\left(\rho_{{}_{0}}+3p_{r_{0}})(\rho_{{}_{2}}+p_{r_{2}}\right)-\frac{32\pi^{2}}{9}\rho_{{}_{0}}(\rho_{{}_{0}}+3p_{r_{0}})(\rho_{{}_{0}}+p_{r_{0}})$
$\hskip-11.38092pt+8\pi\alpha\left(\frac{\rho_{{}_{2}}p_{r_{0}}}{5}+\frac{\rho_{{}_{0}}p_{r_{2}}}{3}\right),$ (B.17)
$\hskip-11.38092pt\Psi_{{}_{4}}=\frac{1}{(\rho_{{}_{0}}+p_{r_{0}})}\left[8\pi\alpha\left(\frac{\rho_{{}_{2}}p_{r_{0}}}{5}+\frac{\rho_{{}_{0}}p_{r_{2}}}{3}\right)-\frac{\Psi_{{}_{2}}}{2}(\rho_{{}_{2}}+p_{r_{2}})-2p_{r_{4}}\right],$ (B.18)
$\hskip-11.38092pt\sigma_{{}_{4}}=\frac{16\pi\alpha}{3}\rho_{{}_{0}}p_{r_{2}}.$ (B.19)
Note that all coefficients shown in Eqs. (B.13)-(B.19) are concord with those obtained for the isotropic case, since when $\alpha=0$, Eqs. (B.14)-(B.15) and (B.17)-(B.18) reproduce the equalities reported in Ref. [19] and with the equalities (B.16) and (B.19) being identically null. Moreover, by evaluating the perturbation equation to the second-order terms $\left[{\tilde{H}}_{1}^{\prime\prime}(0),{\tilde{K}}^{\prime\prime}(0),{\tilde{W}}^{\prime\prime}(0),{\tilde{X}}^{\prime\prime}(0)\right]$, we find that the following relationships must hold:
$\hskip-11.38092pt\frac{\ell+3}{2}{\tilde{H}}_{1}^{\prime\prime}(0)-{\tilde{K}}^{\prime\prime}(0)+8\pi(p_{r_{0}}+\rho_{{}_{0}})\frac{\ell+3}{\ell(\ell+1)}{\tilde{W}}^{\prime\prime}(0)=4\pi\left[\frac{1}{3}(2\ell+3)\rho_{{}_{0}}-p_{r_{0}}\right]{\tilde{H}}_{1}(0)$
$\hskip-11.38092pt+8\pi(p_{r_{0}}+\rho_{{}_{0}})Q_{1}+\frac{1}{2}Q_{0}-\frac{8\pi}{\ell}\left(p_{r_{2}}+\rho_{{}_{2}}+\frac{16\pi\alpha}{3}\rho_{{}_{0}}p_{r_{0}}\right){\tilde{W}}(0),$ (B.20)
$\hskip-11.38092pt\frac{\ell+2}{2}{\tilde{K}}^{\prime\prime}(0)-\frac{\ell(\ell+1)}{4}{\tilde{H}}_{1}^{\prime\prime}(0)-4\pi(p_{r_{0}}+\rho_{{}_{0}}){\tilde{W}}^{\prime\prime}(0)=4\pi\left[p_{r_{2}}+\rho_{{}_{2}}+\frac{8\pi}{3}\rho_{0}(p_{r_{0}}+\rho_{{}_{0}})\right]{\tilde{W}}(0)$
$\hskip-11.38092pt+\frac{1}{2}Q_{0}+\left(\frac{4\pi}{3}\rho_{{}_{0}}+4\pi p_{r_{0}}\right){\tilde{K}}(0),$ (B.21)
$\hskip-11.38092pt\frac{\ell+2}{2}{\tilde{X}}^{\prime\prime}(0)+e^{\Psi_{0}}(p_{r_{0}}+\rho_{{}_{0}})\left[\frac{\omega^{2}}{2}e^{-2\Psi_{{}_{0}}}+2\pi\left(p_{r_{0}}+\rho_{{}_{0}}\right)-\frac{\ell+2}{2}\Psi_{{}_{2}}+\frac{\ell+1}{2}\frac{\sigma_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right]{\tilde{W}}^{\prime\prime}(0)$
$\hskip-11.38092pt+\alpha e^{\Psi_{{}_{0}}}p_{r_{0}}{\tilde{K}}^{\prime\prime}(0)+\frac{\ell(\ell+1)}{8}e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}}){\tilde{H}}_{1}^{\prime\prime}=\left[\frac{\ell}{2}\left[(p_{r_{0}}+\rho_{{}_{0}})\Psi_{{}_{2}}+p_{r_{2}}+\rho_{{}_{2}}+\sigma_{{}_{2}}\right]\right$.
$\hskip-11.38092pt\left.+\frac{d\rho}{dp_{r}}\sigma_{{}_{2}}\right]\frac{{\tilde{X}}(0)}{p_{r_{0}}+\rho_{{}_{0}}}+e^{\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})\biggl{[}\left(-\Psi_{{}_{2}}+\frac{3\sigma_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right){\tilde{K}}(0)-\left(\frac{\alpha p_{r_{0}}}{p_{r_{0}}+\rho_{{}_{0}}}+\frac{1}{4}\right)Q_{0}$
$\hskip-11.38092pt-\frac{\omega^{2}}{2}e^{-2\Psi_{{}_{0}}}{\tilde{H}}_{1}(0)-\frac{\ell(\ell+1)}{2}\left(\Psi_{{}_{2}}-\frac{\sigma_{{}_{2}}}{p_{r_{0}}+\rho_{{}_{0}}}\right)Q_{1}+\biggl{[}(\ell+2)\Psi_{{}_{4}}+\frac{\ell\sigma_{{}_{4}}}{2(p_{r_{0}}+\rho_{{}_{0}})}$
$\hskip-11.38092pt-2\pi\left(p_{r_{2}}+\rho_{{}_{2}}+\sigma_{{}_{2}}\right)-\frac{16\pi^{2}}{3}\rho_{{}_{0}}(p_{r_{0}}+\rho_{{}_{0}})-\frac{4\pi}{3}\rho_{{}_{0}}\Psi_{{}_{2}}+\omega^{2}e^{-2\Psi_{{}_{0}}}\left(\Psi_{{}_{2}}-\frac{4\pi}{3}\rho_{{}_{0}}\right)\biggr{]}{\tilde{W}}(0)\biggr{]}$ (B.22)
$\hskip-11.38092pt-\frac{1}{4}(p_{r_{0}}+\rho_{{}_{0}}){\tilde{K}}^{\prime\prime}(0)-\frac{1}{2}\left[p_{r_{2}}+\frac{\ell+3}{\ell(\ell+1)}\omega^{2}e^{-2\Psi_{{}_{0}}}\left(p_{r_{0}}+\rho_{{}_{0}}\right)\right]{\tilde{W}}^{\prime\prime}(0)-\frac{1}{2}e^{-\Psi_{{}_{0}}}{\tilde{X}}^{\prime\prime}(0)$
$+\frac{\alpha}{2}\left[p_{r_{2}}{\tilde{K}}(0)+p_{r_{0}}\left(Q_{0}+{\tilde{K}}^{\prime\prime}(0)\right)\right]=-\frac{1}{2}e^{-\Psi_{{}_{0}}}\Psi_{{}_{2}}{\tilde{X}}(0)+\frac{1}{4}(p_{r_{2}}+\rho_{{}_{2}}){\tilde{K}}(0)$
$+\frac{1}{4}\left(p_{r_{0}}+\rho_{{}_{0}}\right)Q_{0}+\frac{1}{2}\alpha p_{r_{0}}\Psi_{{}_{2}}{\tilde{K}}(0)+\left[\frac{\omega^{2}}{\ell}e^{-2\Psi_{{}_{0}}}\left[\frac{1}{2}(p_{r_{2}}+\rho_{{}_{2}}+\sigma_{{}_{2}})-(p_{r_{0}}+\rho_{{}_{0}})\Psi_{{}_{2}}\right]\right$.
$\hskip-11.38092pt\left.+p_{r_{4}}-\frac{4\pi}{3}\rho_{{}_{0}}p_{r_{2}}\right]{\tilde{W}}(0)-\frac{1}{2}\omega^{2}e^{-2\Psi_{{}_{0}}}(p_{r_{0}}+\rho_{{}_{0}})Q_{1},$ (B.23)
with $Q_{0}$ and $Q_{1}$ defined by
$\hskip-11.38092ptQ_{0}=\frac{4}{(\ell+2)(\ell-1)}\biggl{[}-8\pi e^{-\Psi_{{}_{0}}}{\tilde{X}}(0)-\left(\omega^{2}e^{-2\Psi_{{}_{0}}}+\frac{8\pi}{3}\rho_{{}_{0}}\right){\tilde{K}}(0)+\left[\omega^{2}e^{-2\Psi_{{}_{0}}}\right$.
$\hskip-11.38092pt\left.-\ell(\ell+1)\left(\frac{2\pi}{3}\rho_{{}_{0}}+2\pi p_{r_{0}}\right)\right]{\tilde{H}}_{1}{(0)}-\frac{128\pi^{2}}{3}\alpha\rho_{{}_{0}}p_{r_{0}}{\tilde{W}}(0)\biggr{]},$ (B.24)
$\hskip-11.38092ptQ_{1}=\frac{2}{\ell(\ell+1)}\left[\frac{4\pi}{3}(\ell+1)\rho_{{}_{0}}{\tilde{W}}(0)-\frac{3}{2}{\tilde{K}}(0)+\frac{e^{-\Psi_{{}_{0}}}}{\Gamma p_{r_{0}}}{\tilde{X}}(0)\right].$ (B.25)
It is important to mention that all equations presented in this section agree with those presented in [40, 19] for the isotropic case.